# An LLM Agent for Functional Bug Detection in Network Protocols

### v1

@@ Abstract
Functional correctness is critical for ensuring the reliability and security of network protocol implementations. Functional bugs, instances where implementations diverge from behaviors specified in RFC documents, can lead to severe consequences, including faulty routing, authentication bypasses, and service disruptions. Detecting these bugs requires deep semantic analysis across specification documents and source code, a task beyond the capabilities of traditional static analysis tools. This paper introduces RFCScan, an autonomous agent that leverages large language models (LLMs) to detect functional bugs by checking conformance between network protocol implementations and their RFC specifications. Inspired by the human auditing procedure, RFCScan comprises two key components: an indexing agent and a detection agent. The former hierarchically summarizes protocol code semantics, generating semantic indexes that enable the detection agent to narrow down the scanning scope. The latter employs demand-driven retrieval to iteratively collect additional relevant data structures and functions, eventually identifying potential inconsistencies with the RFC specifications effectively. We evaluate RFCScan across six real-world network protocol implementations. RFCScan identifies 47 functional bugs with 81.9% precision, of which 20 bugs have been confirmed or fixed by developers.
@@ I Introduction
To enhance the reliability and security of network protocols, it is essential to identify functional bugs in their implementations. Manually detecting such bugs is notoriously difficult, as developers have to compare the implementation against every relevant clause in lengthy RFC documents—a process that demands significant human effort and domain expertise. However, automating functional bug detection remains a challenging task. First, the semantic properties depicting protocol functionality are highly diverse and expressed informally in natural language within RFC documents. In contrast to generic low-level safety properties [3], which can often be uniformly represented using logical constraints, application-specific high-level semantic properties (or semantic properties in this paper), such as those found in network protocols, are considerably more difficult to formalize due to their reliance on domain knowledge and context-dependent behaviors. As shown in the example in Figure 1(a), such properties in the RFC document often require reasoning about the states of multiple protocol entities and the complex interactions among them, making encoding such properties inherently intricate. Second, the diversity of semantic properties significantly hinders the development of effective detection mechanisms. Conventional bug detection techniques, such as constraint-based methods (e.g., symbolic execution [4, 5, 6]) and graph-based reasoning approaches (e.g., data-flow reachability analysis [7, 8]), are inadequate for automatically validating the semantic properties of network protocols. These methods cannot automatically localize the functions relevant to specific properties and verify their compliance with the specification. For instance, it is challenging to localize the function route_lost shown in Figure 1(b) and leverage a uniform reasoning framework to detect violations of the associated semantic property.
@@
Although recent studies have proposed various techniques for bug detection in network protocol implementations, they remain inadequate to effectively identify functional bugs. In general, existing approaches fall into three categories. First, fuzzing techniques aim to uncover bugs by generating crash-triggering inputs [9, 10]. However, functional bugs typically do not cause crashes or other observable anomalies, making them difficult for fuzzing-based methods to detect. Second, differential analysis can reveal behavioral inconsistencies and discover potential functional bugs by comparing multiple independent implementations of the same protocol [11, 12, 13]. Yet, this approach depends on the availability of high-quality alternative implementations and does not guarantee conformance to the original specification, thereby limiting its recall for functional bug detection. Third, formal verification offers strong semantic guarantees but requires rigorous formalization of protocol properties and formal reasoning frameworks [14, 15]. These requirements may not be easily achievable for real-world protocols due to the informal and diverse specifications highlighted above. These limitations suggest the need for a new paradigm of bug detection that enables semantic reasoning on both informal specifications and implementation behavior, an essential capability for effectively detecting functional bugs.
@@
This paper presents RFCScan, the first LLM-based agent for detecting functional bugs in network protocol implementations. Our approach is motivated by the observation that LLMs exhibit strong capabilities in understanding both natural language specifications, such as RFC documents, and the semantics of code snippets when they are well scoped, such as those spanning just a few related functions. Notably, RFC documents are typically well-structured and articulate protocol specifications clearly in natural language. Meanwhile, protocol implementations often follow intuitive naming conventions that reflect the functionality of functions and modules. This suggests a promising opportunity: Given an informal property described in an RFC, we can leverage LLMs to align the natural language description of the property with the code structure. If the functions related to the property can be precisely identified by further code retrieval (e.g., along calling contexts), LLMs can perform comprehensive and accurate reasoning over these functions to discover potential functional inconsistencies, thereby enabling functional bug detection. Technically, RFCScan consists of two collaborative agents, namely an indexing agent and a detection agent, which perform the following core analyses, respectively.
@@
  • Code Semantic Indexing. To bridge informal properties in RFC documents with relevant code scope, the indexing agent constructs hierarchical semantic indexes of the protocol implementation using LLMs, which summarizes the semantics of functions, files, and directories into concise natural language descriptions. Such semantic indexes allow RFCScan to narrow down the scope of functions relevant to the target property, significantly reducing token and time costs during the detection.
@@
  • Retrieval-Guided Detection. To collect sufficient contexts, the detection agent progressively retrieves additional relevant functions using symbolic tools driven by the LLM. If the currently available context suffices to validate or refute conformance to the semantic property, RFCScan proceeds to the next property. Otherwise, it augments the context by retrieving callers or callees until a conclusion can be drawn. The retrieved functions can also support self-critics in the detection, during which the detection agent reviews its own reasoning steps upon the retrieved functions to mitigate hallucinations and improve the precision of the detection.
@@
We implement RFCScan as a prototype, powered by Claude 3.5 Sonnet. We apply RFCScan to the implementations of six widely-used network protocols, including Babel, DHCP, and IGMP, and assess its ability to detect functional bugs according to corresponding RFC documents. It is shown that RFCScan successfully identifies 47 zero-day functional bugs with a high precision of 81.9%. In comparison, existing advanced agents, such as Copilot powered with three advanced models: Claude 3.7 Sonnet, Claude 3.5 Sonnet, and GPT-4o, only identify 26 bugs on average, while exhibiting a substantially higher average false positive rate of 77.5%. We further conduct a comprehensive ablation study, which highlights the significant precision gains attributed to the designs of code semantic indexing and retrieval-guided detection. As of the time of submission, 20 of the detected bugs have been confirmed or fixed by the respective developers. In summary, our work makes the following contributions:
@@
  • We propose RFCScan, the first LLM agent that checks the compliance of network protocol implementation and RFC documents for functional bug detection.
@@
  • We evaluate RFCScan on six network protocol implementations and identify 47 functional bugs with 81.9% precision, outperforming existing LLM-based baselines in both precision and semantic coverage.
  • A prototype implementation of RFCScan is publicly available at https://github.com/zmw12306/RFCScan.
@@ II Overview
In this section, we begin with a real-world functional bug in network protocol and discuss the limitations of existing approaches (Section II-A). Then we provide the key ideas of our approach (Section II-B) with illustrative examples.
@@ II-A Motivating Example
Figure 2: The workflow of RFCScan
@@
Figure 1 shows a real-world functional bug in the implementation of the protocol Babel, a collaborative project under the Linux Foundation with over 400 contributors, powering major open-source large-scale networking platforms like Microsoft’s SONiC and Amazon’s DENT. According to its RFC document shown in Figure 1 (a), when a node loses all feasible routes to a destination but still holds unexpired unfeasible routes, it MUST send a sequence number (seqno) request to trigger route recovery. An unfeasible route means a neighbor recently advertised a path that is not currently usable, but might become valid again. Sending a seqno request prompts that neighbor to re-announce the route, potentially restoring connectivity. However, the original implementation fails to check for unfeasible routes, which can leave the router in a starved state with no way to recover, causing persistent connectivity loss in dynamic networks. As shown in Figure 1 (b), the fix adds an explicit check for unexpired unfeasible routes and issues a seqno request when such a route is found, ensuring RFC compliance and preventing persistent routing failure. This bug has been confirmed and fixed by the developers.
@@
However, detecting functional bugs is extremely challenging. Traditional bug detection methods, such as fuzzing [16, 9, 17] and static analysis [18, 19], primarily focus on identifying low-level issues like memory corruption. Consequently, these approaches are ineffective against non-crashing protocol non-compliance, such as the bug illustrated in Figure 1. Model checking techniques also encounter significant difficulties in detecting this bug, primarily due to their dependence on explicitly specified and formally defined properties. For instance, tools such as CMC [15] and UPPAAL [20] efficiently verify high-level behaviors like state transitions and message sequencing, while struggling to capture diverse and nuanced semantic properties, exemplified in Figure 1(a). Moreover, even when semantic properties can be explicitly specified, model checking requires extensive manual abstraction efforts, which are often labor-intensive and susceptible to errors, constraining their practical applicability. Lastly, recent studies such as ParDiff [11] and ParCleanse [21] target non-crashing parsing bugs by inferring valid packet formats from code or RFCs. While effective for identifying parser-related issues, these approaches fall short in detecting functional bugs that involve more complex semantic properties in components beyond network packet parsers.
II-B RFCScan in a Nutshell
@@
To bridge the gap, we propose RFCScan, an LLM-powered autonomous agent that identifies semantic inconsistencies between source code and RFC documents. Our approach is motivated by the observation that LLMs, having been pre-trained on vast corpora of natural language and programming code, exhibit strong capabilities in understanding both the specification offered by the RFC document and program semantics in the code. For instance, given the semantic property shown in Figure 1(a) and the buggy implementation in Figure 1(b), LLMs can potentially identify the inconsistency between the RFC documents and the corresponding implementation logic. However, applying LLMs for functional bug detection in network protocols is far from trivial, requiring us to resolve two critical technical challenges.
Figure 3: Illustrative examples of the two agents in RFCScan
@@
  • Even after locating a relevant function, LLMs may still struggle to determine whether the implementation satisfies the intended property. Since a semantic requirement is often implemented through multiple interconnected functions spread across files or modules, LLMs typically need additional contexts, such as the functions find_best_route and send_update_resend in Fig. 1.
To tackle these challenges, we draw inspiration from how human developers typically audit functional correctness against protocol specifications. When verifying whether a particular semantic property is correctly implemented, experienced developers first perform a preliminary walkthrough of the codebase to become familiar with the roles of different functions, files, and modules. They then progressively gather additional program constructs, such as related functions or data structures, to reason about the correctness of the implementation and identify potential violations.
RFCScan mirrors this human auditing process through two key technical components: code semantic indexing and retrieval-guided detection. The former creates a semantic map of the codebase by summarizing the functionality of functions, files, and directories in natural language, allowing fast and informed localization of potentially relevant code. The latter incrementally augments the reasoning context with related program constructs, enabling the system to validate a property when its implementation spans multiple functions or modules. These two techniques are realized through two collaborative agents in RFCScan, the indexing agent and the detection agent, whose respective workflows are illustrated in Figure 2.
@@
  • Detection Agent. Given a semantic property, the detection agent uses hierarchical indexes to identify inconsistencies. As shown in Figure 3(b), it proceeds with four actions in five steps: (1) Use the semantic indexes to localize a relevant function, such as route_lost; (2) Detect inconsistencies based on this context; (3) Retrieve additional related functions, such as find_best_route and send_update_resend, if the context is insufficient; (4) Draw a conclusion about a potential violation; and (5) Apply a self-critique strategy to reassess the reasoning for hallucination mitigation.
@@ III Our Approach
RFCScan emulates the reasoning process of human developers when auditing for functional bugs and consists of two collaborative agents, the indexing agent and the detection agent, which support scalable and effective functional bug detection upon network protocol implementations. In what follows, we present more technical details of the two agents with concrete prompts and illustrative examples.
@@ III-A Phase 1: Code Semantic Indexing
To bridge RFC documents with relevant code segments, we follow the behavior of human developers by constructing semantic indexes over the protocol implementation. This approach enables us to capture the semantics of the code in a manner analogous to how developers form a high-level understanding of the implementation. Basically, a simple indexing design is to utilize the names of functions, files, and directories as summaries, as they often suggest the intended functionality [22, 23]. However, such names are not informative enough to capture precise semantics. For example, in the Babel implementation shown in Figure 1, both babeld.c and babel_main.c reside in the same directory, yet their names do not clearly reveal their distinct functionalities. To generate more informative semantic indexes, we leverage LLMs to summarize the contents of functions, files, and directories, producing concise natural language descriptions that form a hierarchical semantic view of the implementation.
As shown in Figure 4, the indexing agent constructs hierarchical semantic indexes in a bottom-up manner. It begins at the function level, where each function definition is passed to the LLM with a prompt, which asks for a concise summary of its semantics. At the file level, summaries of all functions within the same source file are aggregated and used to prompt the LLM for a description of the overall functionality of the file, capturing how individual behaviors contribute to the role of the file in the protocol implementation. This process continues at the directory level, where the agent combines summaries of all files and subdirectories to generate a higher-level description of the directory. Finally, summaries from all directories are aggregated into a repository-level summary, which is a top-level directory summary, reflecting the system’s overall structure and responsibilities. This bottom-up indexing process produces a structured semantic index aligned with the code hierarchy, enabling precise and efficient retrieval during detection. The corresponding prompt templates for each indexing level are shown on the right side of Figure 4. Notably, the code semantic indexing is one-time effort. When the codebase changes, we can incrementally update the indexes. In our evaluation, we choose the Claude 3.5 Sonnet model and the average cost is 1.89\$, shown in Table V. We can also consider using cheaper models for indexing as the task is relatively simple and does not rely on strong reasoning ability.
@@
Consider the repository structure shown on the left side of Figure 3(a). We first ask the LLMs to summarize each function. For example, route_lost handles route loss, find_best_route searches for the best acceptable route, and send_update_resend sends and records update messages. Next, the function-level summaries under the same file are combined to produce a file-level summary that captures the main responsibility of each file. In this case, we can discover that route.c focuses on route maintenance, and message.c on inter-node communication. Finally, all file summaries are merged into a directory-level summary for the babeld module. Lastly, we can obtain the hierarchical semantic indexes shown on the right side of Figure 3(b).

### v2

# RFCAudit: AI Agent for Auditing Protocol Implementations Against RFC Specifications

@@ Abstract
Functional correctness is critical for ensuring the reliability and security of network protocol implementations. Functional bugs, instances where implementations diverge from behaviors specified in RFC documents, can lead to severe consequences, including faulty routing, authentication bypasses, and service disruptions. Detecting these bugs requires deep semantic analysis across specification documents and source code, a task beyond the capabilities of traditional static analysis tools. This paper introduces RFCAudit, an autonomous agent that leverages large language models (LLMs) to detect functional bugs by checking conformance between network protocol implementations and their RFC specifications. Inspired by the human auditing procedure, RFCAudit comprises two key components: an indexing agent and a detection agent. The former hierarchically summarizes protocol code semantics, generating semantic indexes that enable the detection agent to narrow down the scanning scope. The latter employs demand-driven retrieval to iteratively collect additional relevant data structures and functions, eventually identifying potential inconsistencies with the RFC specifications effectively. We evaluate RFCAudit across six real-world network protocol implementations. RFCAudit identifies 47 functional bugs with 81.9% precision, of which 20 bugs have been confirmed or fixed by developers.
@@ I Introduction
To enhance the reliability and security of network protocols, it is essential to identify functional bugs in their implementations. Manually detecting such bugs is notoriously difficult, as developers have to compare the implementation against every relevant clause in lengthy RFC documents—a process that demands significant human effort and domain expertise. However, automating functional bug detection remains a challenging task. First, the semantic properties depicting protocol functionality are highly diverse and expressed informally in natural language within RFC documents. In contrast to generic low-level safety properties [3], which can often be uniformly represented using logical constraints, application-specific high-level semantic properties [4, 5, 6, 7] (or semantic properties in this paper), such as those found in network protocols, are considerably more difficult to formalize due to their reliance on domain knowledge and context-dependent behaviors. As shown in the example in Figure 1(a), such properties in the RFC document often require reasoning about the states of multiple protocol entities and the complex interactions among them, making encoding such properties inherently intricate. Second, the diversity of semantic properties significantly hinders the development of effective detection mechanisms. Conventional bug detection techniques, such as constraint-based methods (e.g., symbolic execution [8, 9, 10, 11, 12]) and graph-based reasoning approaches (e.g., data-flow reachability analysis [13, 14]), are inadequate for automatically validating the semantic properties of network protocols. These methods cannot automatically localize the functions relevant to specific properties and verify their compliance with the specification. For instance, it is challenging to localize the function route_lost shown in Figure 1(b) and leverage a uniform reasoning framework to detect violations of the associated semantic property.
@@
Although recent studies have proposed various techniques for bug detection in network protocol implementations, they remain inadequate to effectively identify functional bugs. In general, existing approaches fall into three categories. First, fuzzing techniques aim to uncover bugs by generating crash-triggering inputs [15, 16, 17]. However, functional bugs typically do not cause crashes or other observable anomalies, making them difficult for fuzzing-based methods to detect. Second, differential analysis can reveal behavioral inconsistencies and discover potential functional bugs by comparing multiple independent implementations of the same protocol [18, 19, 20]. Yet, this approach depends on the availability of high-quality alternative implementations and does not guarantee conformance to the original specification, thereby limiting its recall for functional bug detection. Third, formal verification offers strong semantic guarantees but requires rigorous formalization of protocol properties and formal reasoning frameworks [21, 22]. These requirements may not be easily achievable for real-world protocols due to the informal and diverse specifications highlighted above. These limitations suggest the need for a new paradigm of bug detection that enables semantic reasoning on both informal specifications and implementation behavior, an essential capability for effectively detecting functional bugs.
@@
This paper presents RFCAudit, the first LLM agent for detecting functional bugs in network protocol implementations. Our approach is motivated by the observation that LLMs exhibit strong capabilities in understanding both natural language specifications, such as RFC documents, and the semantics of code snippets when they are well scoped, such as those spanning just a few related functions. Notably, RFC documents are typically well-structured and articulate protocol specifications clearly in natural language. Meanwhile, protocol implementations often follow intuitive naming conventions that reflect the functionality of functions and modules. This suggests a promising opportunity: Given an informal property described in an RFC, we can leverage LLMs to align the natural language description of the property with the code structure. If the functions related to the property can be precisely identified by further code retrieval (e.g., along calling contexts), LLMs can perform comprehensive and accurate reasoning over these functions to discover potential functional inconsistencies, thereby enabling functional bug detection. Technically, RFCAudit consists of two collaborative agents, namely an indexing agent and a detection agent, which perform the following core analyses, respectively.
@@
  • Code Semantic Indexing. To bridge informal properties in RFC documents with relevant code scope, the indexing agent constructs hierarchical semantic indexes of the protocol implementation using LLMs, which summarizes the semantics of functions, files, and directories into concise natural language descriptions. Such semantic indexes allow RFCAudit to narrow down the scope of functions relevant to the target property, significantly reducing token and time costs during the detection.
@@
  • Retrieval-Guided Detection. To collect sufficient contexts, the detection agent progressively retrieves additional relevant functions using symbolic tools driven by the LLM. If the currently available context suffices to validate or refute conformance to the semantic property, RFCAudit proceeds to the next property. Otherwise, it augments the context by retrieving callers or callees until a conclusion can be drawn. The retrieved functions can also support self-critics in the detection, during which the detection agent reviews its own reasoning steps upon the retrieved functions to mitigate hallucinations and improve the precision of the detection.
@@
We implement RFCAudit as a prototype [23], powered by Claude 3.5 Sonnet. We apply RFCAudit to the implementations of six widely-used network protocols, including Babel, DHCP, and IGMP, and assess its ability to detect functional bugs according to the corresponding RFC documents. It is shown that RFCAudit successfully identifies 47 zero-day functional bugs with a high precision of 81.9%. In comparison, existing advanced agents, such as Copilot powered with three advanced models: Claude 3.7 Sonnet, Claude 3.5 Sonnet, and GPT-4o, only identify 26 bugs on average, while exhibiting a substantially higher average false positive rate of 77.5%. We further conduct comprehensive ablation studies, which highlight the significant precision gains attributed to the designs of code semantic indexing and retrieval-guided detection. In summary, our work makes the following contributions:
@@
  • We propose RFCAudit, the first LLM agent that checks the compliance of network protocol implementation and RFC documents for functional bug detection.
@@
  • We evaluate RFCAudit on six network protocol implementations and identify 47 functional bugs with 81.9% precision, outperforming existing LLM-based baselines in both precision and semantic coverage. 20 of the detected bugs have been confirmed or fixed by the respective developers.
@@ II Overview
In this section, we begin with a real-world functional bug in a network protocol implementation and discuss the limitations of existing approaches (Section II-A). Then we provide the key ideas of our approach (Section II-B) with illustrative examples.
@@ II-A Motivating Example
Figure 2: The workflow of RFCAudit
@@
Figure 1 shows a real-world functional bug in FRRouting, a collaborative project under the Linux Foundation with over 400 contributors, powering major open-source large-scale networking platforms like Microsoft’s SONiC and Amazon’s DENT. According to the RFC document shown in Figure 1(a), when a node loses all feasible routes to a destination but still holds unexpired unfeasible routes, it MUST send a sequence number (seqno) request to trigger route recovery. An unfeasible route means a neighbor recently advertised a path that is not currently usable, but might become valid again. Sending a seqno request prompts that neighbor to re-announce the route, potentially restoring connectivity. However, the original implementation fails to check for unfeasible routes, which can leave the router in a starved state with no way to recover, causing persistent connectivity loss in dynamic networks. As shown in Figure 1(b), the fix adds an explicit check for unexpired unfeasible routes and issues a seqno request when such a route is found, ensuring RFC compliance and preventing persistent routing failure. This bug has been confirmed and fixed by the developers.
@@
However, detecting functional bugs is extremely challenging [18, 24]. Traditional bug detection methods, such as fuzzing [25, 15, 26] and static analysis [27, 28], primarily focus on identifying low-level issues like memory corruption. Consequently, these approaches are ineffective against non-crashing protocol non-compliance, such as the bug illustrated in Figure 1. Model checking techniques also encounter significant difficulties in detecting this bug, primarily due to their dependence on explicitly specified and formally defined properties. For instance, tools such as CMC [22] and UPPAAL [29] efficiently verify high-level behaviors like state transitions and message sequencing, while struggling to capture diverse and nuanced semantic properties, exemplified in Figure 1(a). Moreover, even when semantic properties can be explicitly specified, model checking requires extensive manual abstraction efforts, which are often labor-intensive and error-prone, constraining their practical applicability. Lastly, recent studies such as ParDiff [18] and ParCleanse [30] target non-crashing parsing bugs by inferring valid packet formats from code or RFCs. Although effective in identifying parser-related issues, these approaches fall short in detecting functional bugs that involve more complex semantic properties in components beyond network packet parsers.
II-B RFCAudit in a Nutshell
@@
To bridge the gap, we propose RFCAudit, an LLM-powered autonomous agent that identifies semantic inconsistencies between source code and RFC documents. Our approach is motivated by the observation that LLMs, having been pre-trained on vast corpora of natural language and programming code, exhibit strong capabilities in understanding both the specification offered by the RFC document and program semantics in the code. For instance, given the semantic property shown in Figure 1(a) and the buggy implementation in Figure 1(b), LLMs can potentially identify the inconsistency between the RFC documents and the corresponding implementation logic. However, applying LLMs for functional bug detection in network protocols is far from trivial, requiring us to resolve two critical technical challenges.
Figure 3: Illustrative examples of the two agents in RFCAudit
@@
  • Even after locating a relevant function, LLMs may still struggle to determine whether the implementation satisfies the intended property. Since a semantic requirement is often implemented through multiple interconnected functions spread across files or modules, LLMs typically need additional contexts, such as the functions find_best_route and send_update_resend in Figure 1.
To tackle these challenges, we draw inspiration from the way human developers typically audit functional correctness against protocol specifications. When verifying whether a particular semantic property is correctly implemented, experienced developers first perform a preliminary walkthrough of the codebase to become familiar with the roles of different functions, files, and modules. They then progressively gather additional program constructs, such as related functions or data structures, to reason about the correctness of the implementation and identify potential violations.
RFCAudit mirrors this human auditing process through two key technical components: code semantic indexing and retrieval-guided detection. The former creates a semantic map of the codebase by summarizing the functionality of functions, files, and directories in natural language, allowing fast and informed localization of potentially relevant code. The latter incrementally augments the reasoning context with related program constructs, enabling the system to validate a property when its implementation spans multiple functions or modules. These two techniques are realized through two collaborative agents in RFCAudit, the indexing agent and the detection agent, whose respective workflows are illustrated in Figure 2.
@@
  • Detection Agent. Given a semantic property, the detection agent uses hierarchical indexes to identify inconsistencies. As shown in Figure 3(b), it proceeds with four actions in five steps: (1) Use the semantic indexes to localize a relevant function, such as route_lost; (2) Detect inconsistencies based on this context; (3) Retrieve additional relevant functions, such as find_best_route and send_update_resend, if the context is insufficient; (4) Draw a conclusion about a potential violation; and (5) Apply a self-critique strategy to reassess the reasoning for hallucination mitigation.
@@ III Our Approach
RFCAudit emulates the reasoning process of human developers when auditing for functional bugs and consists of two collaborative agents, the indexing agent and the detection agent, which support scalable and effective functional bug detection upon network protocol implementations. In what follows, we present more technical details of the two agents with concrete prompts and illustrative examples.
@@ III-A Phase 1: Code Semantic Indexing
To bridge RFC documents with relevant code segments, we follow the behavior of human developers by constructing semantic indexes over the protocol implementation. This approach enables us to capture the semantics of the code in a manner analogous to the way developers form a high-level understanding of the implementation. Basically, a simple indexing design is to utilize the names of functions, files, and directories as summaries, as they often suggest the intended functionality [31, 32]. However, such names are not always informative enough to capture precise semantics. For example, in the Babel implementation shown in Figure 1, both babeld.c and babel_main.c reside in the same directory, yet their names do not clearly reveal their distinct functionalities. To generate more informative semantic indexes, we leverage LLMs to summarize the contents of functions, files, and directories, producing concise natural language descriptions that form a hierarchical semantic view of the implementation.
As shown in Figure 4, the indexing agent constructs hierarchical semantic indexes in a bottom-up manner. It begins at the function level, where each function definition is passed to the LLM with a prompt, which asks for a concise summary of its semantics. At the file level, summaries of all functions within the same source file are aggregated to prompt the LLM to generate a description of the file’s overall functionality, capturing how individual behaviors contribute to the role of the file in the protocol implementation. This process continues at the directory level, where the agent combines summaries of all files and subdirectories to generate a higher-level description of the directory. Finally, summaries from all directories are aggregated into a repository-level summary, which is a top-level directory summary, reflecting the system’s overall structure and responsibilities. This bottom-up indexing process produces a structured semantic index aligned with the code hierarchy, enabling precise and efficient retrieval during detection. The corresponding prompt templates for each indexing level are shown on the right of Figure 4. Notably, the code semantic indexing is one-time effort. When the codebase changes, we can incrementally update the indexes. In our evaluation, we choose the Claude 3.5 Sonnet model and the average cost for a full repository is \$1.89, shown in Table V. We can also consider using cheaper models for indexing as the task is relatively simple and does not rely on strong reasoning ability.
@@
Consider the repository structure shown on the left of Figure 3(a). We first ask the LLMs to summarize each function. For example, route_lost handles route loss, find_best_route searches for the best acceptable route, and send_update_resend sends and records update messages. Next, the function-level summaries under the same file are combined to produce a file-level summary that captures the main responsibility of each file. In this case, we can discover that route.c focuses on route maintenance, and message.c on inter-node communication. Finally, all file summaries are merged into a directory-level summary for the babeld module. Lastly, we can obtain the hierarchical semantic indexes shown on the right of Figure 3(b).